A. Find the length of the confidence interval (s = standard deviation)




1. s = 3




n = 250




Confidence level = 95%



2. s = 6




n = 400




Confidence level = 99%



B. Determine the sample size, given the following data.




1. s = 5




E = 2.42




Confidence level = 95%



2. You want to estimate the mean gasoline price within your town to the margin of error of 6 centavos. Local newspaper reports the standard deviation for gas price in the area is 30 centavos. What sample size is needed to estimate the mean gas prices at 99% confidence level?



3. Carlos wants to replicate a study where the highest observed value is 14.8 while the lowest is 14.2. He wants to estimate the population mean µ to the margin of error of 0.025 of its true value. Using 95% confidence level, find the sample size n that he need.

1
Expert's answer
2022-05-11T15:33:02-0400

A.

1. The critical value for "\alpha = 0.05" and "df=n-1=249" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 1.969537."


"Length=2t_c\times\dfrac{s}{\sqrt{n}}=2(1.969537)\times\dfrac{3}{\sqrt{250}}"

"=0.7474"

2. The critical value for "\alpha = 0.01" and "df=n-1=399" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.588207."


"Length=2t_c\times\dfrac{s}{\sqrt{n}}=2(2.588207)\times\dfrac{6}{\sqrt{400}}"

"=0.7765"


B.

1.


"E=t_c\times\dfrac{s}{\sqrt{n}}"

The critical value for "\alpha = 0.05" and "df=n-1=18" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.100922."


"E=2.100922\times\dfrac{5}{\sqrt{19}}=2.41"

The critical value for "\alpha = 0.05" and "df=n-1=17" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.1098162."


"E= 2.109816\times\dfrac{5}{\sqrt{18}}=2.49"

"n=19"


2. The critical value for "\alpha = 0.01" is "z_c = z_{1-\alpha/2} = 2.5758."


"E=z_c\times\dfrac{\sigma}{\sqrt{n}}"

"n\ge(\dfrac{z_c\times\sigma}{E})^2"

"n\ge(\dfrac{2.5758\times30}{6})^2"

"n=166"

3.


"\bar{x}=\dfrac{14.2+14.8}{2}=14.5"

"E=0.025(14.5)=0.3625"

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."


"E=z_c\times\dfrac{\sigma}{\sqrt{n}}"

"n\ge(\dfrac{z_c\times\sigma}{E})^2"

"n\ge(\dfrac{1.96\times\sigma}{0.3625})^2"

Let "\sigma=1"

"n=30"